If the ratio of the ages of Kissi and Esinam is 3:5 and that of Esinam and Lariba is 3:5 and
the sum of the ages of all 3 is 147 years, what is the age difference between oldest the
youngest?
1
Expert's answer
2020-06-15T04:41:42-0400

We call Kissi's age X, Esinam's age Y, and Labira's age Z


Step 1) X / Y = 3/5, Y / Z = 3/5.


Step 2) X = Y * 3/5, Z = 5 * Y / 3.


Step 3) X + Y + Z = 147, Y * 3/5 + Y + 5 * Y / 3 = 147, hence Y = 45, X = 27, Z = 75


Step 4) Z-X = 48. Answer 48.